OpenStudy (anonymous):

how do you solve for the complex fraction: (3/2+x)-(4/2-x) /(1/x+2)-(3/x-2)

OpenStudy (phi):

find a common denominator for the tops, and subtract ditto for the bottom fractions then change the division by a fraction by inverting the bottom fraction and then multiplying

OpenStudy (anonymous):

multiply top and bottom by \((x+2)(x-2)\) cancelling merrily as you go clears the compound fraction in one step

OpenStudy (anonymous):

alright i did that and cancelled out everything and got 2 for the answer...is that right?

OpenStudy (phi):

is the problem \[ \frac{ \frac{3}{2+x}-\frac{4}{2-x}}{\frac{1}{x+2}-\frac{3}{x-2}}\]

OpenStudy (anonymous):

yup